10000e^(-0.06t) = 1000
Please help me solve for t :D

Answer by scott8148(6628)

dividing by 10000 ___ e^(-0.06t) = 1/10

rules for exponents ___ e^(.06t) = 10

taking natural log ___ .06t = ln(10)

dividing by .06 ___ t = [ln(10)] / .06